### v1

# across languages

Abstract
As large language models (LLMs) are deployed globally, it is crucial that their responses are calibrated across languages to accurately convey uncertainty and limitations. Previous work has shown that LLMs are linguistically overconfident in English, leading users to overrely on confident generations. However, the usage and interpretation of epistemic markers (e.g., ‘It’s definitely’, ‘I think’) can differ sharply across languages. Here, we study the risks of multilingual linguistic (mis)calibration, overconfidence, and overreliance across five languages to evaluate the safety of LLMs in a global context.
We find that overreliance risks are high across all languages. We first analyze the distribution of LLM-generated epistemic markers, and observe that while LLMs are cross-linguistically overconfident, they are also sensitive to documented linguistic variation. For example, models generate the most markers of uncertainty in Japanese and the most markers of certainty in German and Mandarin. We then measure human reliance rates across languages, finding that while users strongly rely on confident LLM generations in all languages, reliance behaviors differ cross-linguistically: for example, users rely significantly more on expressions of uncertainty in Japanese than in English. Taken together, these results indicate high risk of reliance on overconfident model generations across languages. Our findings highlight the challenges of multilingual linguistic calibration and stress the importance of culturally and linguistically contextualized model safety evaluations.
1 Introduction
A critical component of safe and reliable human-AI interaction is the ability for agents to clearly express their epistemic states, meta-information about their certainty in the knowledge they communicate. One common approach is to have a large language model (LLM) linguistically express its confidence using epistemic markers like ‘It’s definitely’ or ’I think’ (Kadavath et al., 2022; Tian et al., 2023; Liu et al., 2023; Xiong et al., 2024; Tanneru et al., 2023; Mielke et al., 2022; Lin et al., 2022; Stengel-Eskin et al., 2024). Recent work (e.g. Zhou et al., 2024b) has emphasized the need to calibrate model confidence with human reliance. In other words, models should express confidence in a way that triggers the right human reliance behaviors.
Past work has found that English language models are systematically overconfident—that is, they generate epistemic markers expressing high certainty even when incorrect, with high frequency. Compounding this, English speakers tend to systematically overrely on language model outputs (Zhou et al., 2024b). Yet, in the linguistics literature, it is well-established that the (1) usage and (2) interpretation of epistemic markers differ sharply and richly across languages (Davidson & Chrisman, 1994; Itani, 1995; Doupnik & Richter, 2003; Yang, 2013). Here, we contend that the current literature’s focus on English epistemic markers is insufficient for understanding the unique safety harms multilingual models pose to a global audience. Our contributions are as follows:
  1. 1.
    Model overconfidence differs across languages. Although it is well-documented that LLMs generate linguistically confident but incorrect answers in English, it is unclear whether this effect holds across languages. Examining five languages (English, French, German, Japanese, and Mandarin), we find that LLMs are sensitive to distributional differences in the usage of epistemic markers between languages (e.g. more hedges in Japanese than in English), but that overconfidence rates are high across all five languages.
  2. 2.
    Humans reliance differs across languages. Thus far, reliance have focused on English-speaking users (Zhou et al., 2024c). Yet in reality, reliance on epistemic markers will vary based on linguistic and cultural norms. In a bilingual human reliance study, we show that, while high in all languages, reliance rates do differ significantly between languages.
  3. 3.
    Yet overreliance risks remains high across languages. Given our first result, we might naïvely expect that overreliance risks—i.e. relying on overconfident generations—also differ across languages. For example, since LLMs produce more hedges in Japanese than in English, Japanese speakers might actually be less prone to overreliance risk. But taking into account differences in reliance rates, we find that this reasoning is incorrect. Instead, the same bilingual speaker will systematically rely more on Japanese uncertainty markers than their translated English counterparts. In other words, the distribution of human reliance differs across languages in a way that cancels out the effect of distribution shift in LLM overconfidence.
The rest of the paper is structured as follows. In Section 3, we measure whether or not language models are linguistically calibrated in four new languages (French, German, Japanese, and Mandarin). Next (Section 3.3), we analyze if the produced epistemic markers adhere to documented linguistic norms (that is, do LLMs produce more uncertainty markers in languages that are known to hedge more?). Finally, in Section 4, we consider the risk of model miscalibration in terms of human perception of these markers across languages, using Zhou et al. (2024b)’s human reliance evaluation framework.
Taken together, our results implicate significant global safety risks for overreliance. We argue that the appropriate use of epistemic markers necessitates contextualization of expressions to the language of use. While our work here concentrates only on difference in language, we expect that similar contextualization is necessary for differences in topic and setting, as well as other differences in user demographics (e.g. age, gender). Our results highlight the existing risks of LLM overconfidence and the linguistic and social norms that must be considered when developing safe and calibrated language models.
2 Related Work
2.1 LLMs and Uncertainty Expressions
Recent works in NLP have been interested in aligning internal model probabilities with their task accuracy (Jiang et al., 2021; Desai & Durrett, 2020; Jagannatha & Yu, 2020; Kamath et al., 2020; Kong et al., 2020; Hofweber et al., 2024). Most recently, there has been growing interest in calibrating models in a manner more accessible to end users by directly emitting model probabilities as a proxy for model confidence to the user (Kadavath et al., 2022; Tian et al., 2023; Liu et al., 2023; Xiong et al., 2024; Tanneru et al., 2023; Mielke et al., 2022; Lin et al., 2022; Stengel-Eskin et al., 2024). A key finding of this line of work is that English LLMs are miscalibrated to their accuracies: models are less accurate when they produce markers of certainty. Zhou et al. (2023) posit that this is an artifact of (1) pretraining data, which contains a large number of Q/A pairs from online forums in which questions are highly confident and answers often include hedges, and (2) the preference optimization process, as human users ‘prefer’ more certain responses.
Yet, while it is critical to measure whether model confidence is calibrated to generation quality, this is often insufficient to understand the risks of human overreliance. Recent work has instead stressed the need to consider how model calibration affects human behavior. Dhuliawala et al. (2023) and Zhou et al. (2024a) propose alternative methods to measure the risks of overconfident models by directly measuring human reliance via self-incentivized games with AI agents. Studies focused on human reliance and trust rather than language quality alone have provided insights into the features that shape human reliance, such as increased reliance with anthropomorphism (Zhou et al., 2024b) and explanations (Kim et al., 2025), increased trust with medium uncertainty values (Xu et al., 2025).
2.2 Cross-Linguistic Variation
However, a key limitation of existing work on uncertainty expressions in LLMs is its sole focus on model overconfidence in English and among English-speaking participants. Meanwhile, it is well studied in the linguistics literature that epistemic markers vary drastically in form and function across languages and dialects, and these differences are unknown in our understanding of model overconfidence.
Previous work has found that while English and French speakers tend to use mostly moderate expressions of certainty—‘I’m pretty sure,’ ‘I believe’—speakers of Japanese tend to use weakeners—‘I think,’ ‘maybe’—as a baseline (Davidson & Chrisman, 1994; Itani, 1995; Lauwereyns, 2000; Itakura, 2013; Barotto, 2018; Yin et al., 2024). On the other hand, speakers of German and Mandarin tend to use distributions tending on the side of strengtheners (Doupnik & Richter, 2003; Kranich, 2011; Yang, 2013). These effects are robust across speaker identity and domain. [^1] Similar results have been shown for other languages, including Hindi, Turkish, and Italian (Uysal, 2014; Mazzuca et al., 2024). Here, we concentrate on English, French, German, Japanese, and Mandarin as these distributional effects are well-documented and robust, and these languages are well-supported by multilingual LLMs.
Taking our current understanding of LLM overconfidence, this work seems to suggest new miscalibration risks for non-English languages. For one, it may be the case that overconfidence rates differ between languages—for example, if LLMs produce more strengtheners in German or Mandarin, then perhaps they are more overconfident in these languages (and vice versa in Japanese). Further, cross-linguistic distributional differences may lead to differences in reliance patterns between languages. For example, since weakeners are a priori more frequent in Japanese than in English, a user might correspondingly rely more heavily on a weakener in Japanese. In other words, their ‘internal model’ of the LLM might change between languages (in the sense of Gricean recursive reasoning; see Frank & Goodman, 2012; Goodman & Frank, 2016; Degen, 2023).
3 LLM Generations Are Overconfident Across Languages
Past work has found that English LLMs are prone to risks stemming from overconfidence—that is, expressing epistemic certainty even when providing incorrect knowledge. In this section, we investigate the global risks of LLMs by assessing their rate of overconfidence in multiple different languages. We show that multilingual LLMs demonstrate the same overconfidence effects in languages other than English and that LLMs are sensitive to variation in the linguistic norms of using uncertainty in different languages.
A Note On LLM Epistemic States
Previous work has suggested that LLMs rarely produce verbalized expressions of confidence, unless explicitly prompted to do so (e.g. Zhou et al., 2024c). Yet, humans tend to rely on ‘plain’ model generations (i.e. those not containing epistemic markers) even more than they do on strengtheners (Zhou et al., 2024b, also see Figure 5). In other words, unprompted models lead to high risks of overreliance. But this can be mitigated: users tend to be sensitive to linguistic markers of uncertainty (i.e. they rely less on a model when it hedges). As such, here we evaluate models by explicitly eliciting epistemic markers as part of their responses using few-shot prompting.
3.1 Methods
At a high-level, we study the distribution in epistemic markers of model responses to questions from the Massive Multitask Language Understanding benchmark (MMLU; Hendrycks et al., 2020), a multiple-choice QA dataset across 57 subjects, prompting the model to generate epistemic markers as part of its response.
As very few models expressly support multilingual use, we were limited in the breadth of models that could be studied. We analyzed three publicly deployed LLMs at varying scales: GPT-4o (May 2024; Hurst et al., 2024) and Llama 3.1 8B/70B Instruct (July 2024; Grattafiori et al., 2024). [^2] GPT-4o responses were collected in September 2024 for English, German, French, and Japanese, and February 2025 for Mandarin. Note that while all three of these models are designed to be multilingual, the Llama models do not explicitly support Japanese or Mandarin (they do, however, support French and German). As a high-level overview, we first prompted each model to answer multiple-choice questions and to include an epistemic marker in their response. We then annotated these responses by their expressed certainty.
Dataset
We prompted models to respond to multiple-choice questions from a subset of the MMLU test set. Since MMLU is an English-based dataset, we used a parallel, translated version for each target language studied. For Japanese, we used JMMLU, a machine-translated and hand-checked subset [^3] In particular, JMMLU does not include questions from the high_school_government_and_politics, high_school_us_history, and us_foreign_policy categories of the English MMLU. of MMLU. We then machine translated this subset of MMLU from English into the other target languages (French, German, and Mandarin) using the Google Translate API. In total, this resulted in a parallel set of 7,494 questions in each language.
Prompting
In our few-shot prompts, each example response including an expression of uncertainty in addition to the multiple-choice answer. Each prompt included 10 examples, 5 using expressions of uncertainty and 5 using expressions of certainty, in random order.
To generate prompts, for each language we solicited 10 crowdworkers on Prolific to each generate 5 expressions of uncertainty and 5 expressions of certainty. We then selected the top 10 most frequently generated expressions to be used in few-shot examples.
For each prompt, we randomly selected 10 questions from the parallelized MMLU test set as the examples for few-shot. We used the same examples across all languages. We then repeated this process three times, resulting in three sets of few-shot prompts (note that we used the same set of 10 expressions for all three prompts). See Appendix C for an example prompt. In total, we therefore obtained 22,482 generations for each language. Baseline MMLU accuracy by language and model is reported in Table 1. We also report failure rates (i.e. the proportion of times models failed to follow the template and generate epistemic markers) in Table 5; in what follows, we remove failures from all reported results.
Model English French German Japanese Mandarin
GPT-4o 80.96 76.32 77.83 77.08 75.75
Llama-3.1-70B 66.69 57.51 34.47 45.90 55.65
Llama-3.1-8B 59.44 37.17 23.05 31.29 42.17
Table 1: Average MMLU test set accuracy (10-shot, across three trials) across languages by model. Note that across all models, we see consistently higher performance in English. Also observe that while Llama-3.1 does not explicitly support Japanese or Mandarin, both Llama models perform better in Japanese and Mandarin than in German (which is explicitly supported).
Annotation
We then recruited crowdworkers on Prolific to qualitatively annotate the ‘ground truth’ certainty of these model outputs into three categories: weak certainty, moderate certainty, and strong certainty. Under this schema, weak expressions correspond to a high degree of uncertainty, e.g. expressions like ‘I think,’ strong expressions correspond to (near) absolute certainty, e.g. ‘I am 100% sure it is,’ and moderate expressions are in between, e.g. ‘It is most likely.’ We used crowdworker-annotations only for generations from GPT-4o; we trained a classifier on these annotations to label Llama-3.1 generations (see below). Details on annotation are available in Appendix A.
Language Inter-rater reliability ($\kappa$)
French 0.44
English 0.45
German 0.38
Japanese 0.37
Mandarin 0.26444Note we see slightly lower inter-rater reliability in Mandarin compared to other languages. We suspect that this is due to the fact that Mandarin has very few weakeners. In general, we find that raters disagree most between the ‘moderate’ and ‘strong’ categories, and tend to uniformly agree on weakeners; the lack of weakeners thus likely contributes to Mandarin’s lower inter-rater reliability.
Table 2: Inter-rater reliability (Fleiss’ $\kappa$) by language for GPT-4o generations, across three annotators per language. We find fair-to-moderate agreement across annotators in all languages, indicating that annotations are well-calibrated to our universal baseline.
We used three annotators per language, and computed inter-rater reliability using Fleiss’ $\kappa$. We find moderate agreement between annotators for all languages (Table 2), meaning that annotations are well-calibrated to our baseline and thus can be compared across languages.
Classifier
To replicate this process across other models at scale, given budget constraints, we trained a classifier on these annotations for each language. In doing so, we also create an artifact for the community to use in future work on uncertainty requiring annotations.
In each language, we fine-tuned a unique Llama-3.1-8B instance corresponding to each human annotator. [^5] We train for 7 epochs. We used learning rate $5\times 10^{-5}$ and batch size 16. Models were trained on 3-class classification against the labels from each annotator’s GPT-4o generations. We then took the majority vote across all three fine-tuned classifiers, as we did with the human annotators. See Appendix Table 4 for classifier performance; on average, classifiers achieved 78.13% accuracy on a held-out test set. [^6] Each classifier serves as a proxy for a single human annotator. We benchmarked the ‘expected’ accuracy of a classifier by measuring the average accuracy of each human annotator as a classifier; across languages, we found that humans scored 8̃4.14% accuracy on average (see Appendix Table 4). As such, our classifiers perform just slightly worse than a human annotator.
3.2 Results: LLM Calibration and Overconfidence
A key question is whether model generations are calibrated to their true accuracy, or whether they are over/underconfident. To this end, we analyze the relationship between expressed certainty and accuracy on our parallelized MMLU test set.
We measure the overconfidence rate of a model similarly to Zhou et al. (2024c), as the fraction of all responses in which the model uses a strong expression yet gave the wrong answer, i.e. $p(\text{incorrect}\mid\text{strong})$. We find that in all languages, model overconfidence is high: in GPT-4o, the highest performing model (in terms of MMLU accuracy), 15.22% of generations containing strong epistemic markers are incorrect across languages. In Llama 70B and 8B, we similarly find high overconfidence rates of 39.15% and 49.04%. We also find that, in all languages, model responses are at best only somewhat calibrated to accuracy: GPT-4o is slightly more accurate when using strengtheners (as compared to weakeners/moderate markers), while both Llama models show no significant difference across categories.
To examine whether overconfidence risks are worse in non-English languages, we perform this analysis by language (Figure 1 and Table 9). Our findings show that the rate of overconfidence is uniformly higher for non-English languages, with Mandarin having the highest overconfidence rates (18% for GPT-4o, compared to 11% in English). Given that we coded strong expressions as those that convey absolute (or near-absolute) certainty, it is therefore concerning that 15% of these responses are incorrect.
Figure 1: MMLU accuracy by type of epistemic markers. Error bars represent bootstrapped 95% CI. GPT-4o is most calibrated to accuracy, as the model is slightly more accurate when it uses strengtheners. However, in Llama-3.1 models, accuracy is roughly equivalent between all categories of epistemic strength, indicating low calibration.
3.3 LLM Generations Adhere to Linguistic Norms
We are also interested in the baseline distribution of generations across languages—in particular, whether this matches documented linguistic norms. We find that models are generally sensitive to cross-linguistic variation in usage of uncertainty expressions (Figure 2). For example, while in English and French we find that most generations include moderate or strong epistemic markers (96% and 91% in GPT-4o, respectively), we see that in Japanese, most epistemic markers are weakeners (59% in GPT-4o). On the other hand, in German and Mandarin, strengtheners are more prevalent than weak or moderate markers (53% and 80% in GPT-4o, compared to 42% in English).
Figure 2: Distribution of epistemic markers after few-shot prompting. Error bars represent bootstrapped 95% CI. Generations broadly match linguistic norms from the literature, i.e., in Japanese, models produce more hedges (weakeners) than boosters (strengthers), the opposite effect appears in German and Mandarin.
4 Humans Overrely on LM Generations Across Languages
Our previous sections highlight that LLM overconfidence is a risk across all languages; we are curious to know how native speakers of these languages are responsive to variation in the strength of epistemic markers used by LLMs. We approach this question using Zhou et al. (2024b)’s Rel-A.I. framework, creating a task in which we measure human reliance on LLM generations of varying levels of certainty. The motivation behind this work is to measure safety in terms of human behaviors, rather than just what is produced in generations.
We find that users overrely on generations including strengtheners in all languages—that is, they systematically rely on strengtheners at a rate leading to high risk of relying on incorrect generations. Further we also find cross-linguistic variation in the distribution of reliance.
4.1 Methods
Task Setup
We are interested in how the same human might rely on LLMs differently when encountering epistemic markers in different languages. To achieve this, we recruit bilingual speakers and have the same human complete a bilingual task where we then measure human reliance in the target language compared to reliance in English.
The task used 60 trivia questions from Zhou et al. (2024c), which consisted of the 60 most difficult geography questions on Sporcle, an online trivia platform. Using difficult questions encourages users to rely on the model response rather than their own knowledge. We then machine-translated these questions into each target language with the Google Translate API.
During the task, participants were told they would be interacting with an AI agent to answer trivia questions. Each item consisted of a question (e.g. What is the capital of Kiribati?) and the beginning of a model response (e.g. I think it’s…). They were then asked whether they would choose to rely on the model response or if they would rather look up the answer themselves. See Appendix Figure 4 for an example.
Participants were shown 30 question/response pairs in English, and 30 in the target language. Questions were randomly sampled from the set of 60 for each participant, such that the English and target language sets varied between participants. Language order was randomized for each participant (i.e. English vs. the target language), and item order was randomized within each language (i.e. participants saw all generations in one language followed by all generations in the second), mitigating exposure/learning biases that could be introduced from order.
Response Selection
For each language, we select 25 model generations from the previous task. We use 5 generations that were annotated as strong markers, 5 weak markers, and 15 moderate markers. We also include 5 plain expressions for each language, e.g. ‘The answer is…’, which we generated by hand. These plain expressions act as a control to benchmark human reliance when LLMs do not use epistemic markers.
Participants
For each language, we recruited 45 bilingual participants on Prolific, excluding participants from the previous generation and annotation tasks. In this study, we focus on English, German, French, and Japanese. [^7] As the models generally failed to generate weak epistemic markers in Mandarin, we excluded Mandarin for the reliance studies, as it would have resulted in a 2-way instead of 3-way classification task. We filtered participants who self-reported less than C1 (advanced) proficiency.
en fr de jp
Reliance (%) Strengtheners 66.34 54.54 61.90 78.91
Overconfidence Rate GPT-4o 11.26 15.97 15.42 14.85
Llama-3.1-70B 35.63 42.83 66.39 53.14
Llama-3.1-8B 37.98 62.68 76.79 69.64
Overreliance Risk = Reliance $\times$ Overconfidence GPT-4o 7.47 8.71 9.55 11.72
Llama-3.1-70B 23.64 23.36 41.10 41.93
Llama-3.1-8B 25.20 34.19 47.53 54.95
Table 3: Overconfidence rates on MMLU questions, human reliance rates on strengtheners, and expected overreliance risks. Overconfidence rate is defined as the proportion of responses using a strengthener in which the model gives an incorrect answer, $p(\text{incorrect}\mid\text{strong})$. The overreliance risk of a set of generations is the probability of human reliance on a strengthener multiplied by the overconfidence rate. Across all model generations, there is high overreliance risk, with Japanese generations having the highest risk—nearly 1.6 times that of English generations. [^9] English reliance rates were computed by averaging across all bilingual studies.
4.2 Overreliance Across Languages
Our findings illustrate that humans rely heavily on strengtheners across all languages, with an average reliance rate of 65.42% on classified strengtheners (see Table 9 and Figure 5). Although these reliance rates already showcase a concerning risk, these rates as is are insufficient measures of overconfidence risk. What we need instead is a measurement that combines both the rate of overconfidence generations with the rate of overreliance on confident generations. We therefore define the overreliance risk for each model/language as the product of the strengthener reliance rate and the overconfidence rate from Figure 1. This measures the probability that a human will rely on an incorrect response generated by a model using a strengthener.
Using this new metric, we find that overreliance risk is high across languages in models. Even in GPT-4o, the highest accuracy model evaluated, we find an average overreliance risk of nearly 10% across languages. The overreliance risk is substantially worse in Llama-3.1-8B, where users are predicted to rely on incorrect responses 40% of the time.
4.3 Reliance Differences Between Languages
Given that overreliance appears across languages, we might naively assume that, as LLMs produce more hedges in Japanese, Japanese users might face lower risks of overreliance. Yet, we find that this is incorrect: human reliance rates actually differ by category across languages.
The bilingual task allows us to explicitly probe differences in reliance rates between English and the target language, as in Figure 3. We see that German and French do not differ significantly in reliance from English for weak or moderate epistemic markers, and we notice a slight decrease in reliance for plain markers in both languages (Figure 2).
Yet in Japanese, we observe that when the model uses any form of epistemic marker, human reliance increases in comparison to English. This leads to an increased risk of overreliance, as even though models produce more hedges in Japanese, speakers of Japanese are still likely to rely on hedges moreso than in English. In other words, the shift in the distribution of epistemic markers does not actually mitigate overreliance risks. We find that this manifests in the overreliance risks in Table 9: Japanese users are 1.5x more susceptible to overreliance risk than English users.
Figure 3: Differences in human reliance between English and each target language, by condition. Error bars represent bootstrapped 95% CI. Reliance rates on plain expressions is lower in German and French than in English. In German, bilingual users rely less on strengtheners than they do in English. Contrastingly, in Japanese, users rely more on all epistemic markers. French epistemic markers are perceived roughly as they are in English.
5 Discussion and Conclusion
Here, we studied the risks arising from miscalibrated multilingual language models. We found that, while multilingual LLMs do adhere to documented linguistic norms around the production of epistemic markers, they are still systematically overconfident across languages. Further, we showed that users tend to overrely on LLMs in all languages, and that overreliance risks may actually be worse in languages like Japanese, where expressions uncertainty are more common but have diminished function as markers of epistemic state.
More Uncertainty Generations, Less Perceived Uncertainty Our findings illustrate that although languages might differ in their generations of epistemic markers, this does not guarantee a reduction in LLM overreliance. For example, in Japanese generations we see a simultaneous increase in uncertainty markers along with a decrease in the perceived function of hedges—in other words, even though models generate more hedges in Japanese, users are simultaneously more likely to rely on generations containing them. Applying our understanding of how English-speaking users might rely on English markers would have provided us with an inaccurate estimation of overreliance risks of Japanese generations. Our findings show that reliance on the epistemic markers is dependent on the linguistic norms and emphasizes the need to evaluate overreliance in context.
Risks of Emergent Overconfidence Behaviors One of the key advantages of training multilingual LLMs is the transfer learning that happens across languages (Wang & Zheng, 2015). Rather than supplying general knowledge for each individual language, effective transfer learning allows for correct prompt completion in target languages which lack appropriate training data, as long as the data is available in another language (Pires et al., 2019; Ebrahimi & Kann, 2021). However, our work reveals a potential danger of transfer learning in multilingual models. As shown in prior work that language models trained with English might gain an ‘English accent’ (Papadimitriou et al., 2023), our work shares the concern that linguistic norms in English (i.e., overconfidence) might also bias generations in other languages. As LLM training data, annotations, and representations are skewed towards English and Western perspectives (Grattafiori et al., 2024; Hurst et al., 2024; Wendler et al., 2024; Schut et al., 2025), future work ought to ask how these uses of overconfidence could emerge in multilingual models.
Taken together, our results highlight the importance of culturally contextualized and user-centered model safety evaluations.
Acknowledgments
This work benefited from discussion with Zouberou Sayibou, Tatsunori Hashimoto, and other members of the Stanford NLP group. Compute funding was generously supported by Microsoft Accelerating Foundation Models Research and IBM Research.
Ethics Statement
Our work has implications for the design of more fair multilingual language models, and stresses the need for contextualized model safety evaluations. Notably, our analysis targeted only high resource languages, as these were the most available for crowdworker annotations. However, overconfidence and overreliance risks will likely also arise—and may differ—in low-resource languages. Future work should evaluate both the risks we identified here along with potential risks unique to low-resource languages and speaker communities.
All human experiments followed standard IRB protocol and used consent forms to inform the participants of the risks/benefits of the tasks. Crowdworkers were paid \$16 USD per hour. When participants took longer than the expected time we provided bonuses to meet this minimum.
Appendix A Annotation Details
For each language, we first ran a qualifying study on Prolific in which we asked 15 participants—who self-reported C1/C2 proficiency—to annotate 15 model generations on a 1-7 Likert scale of their certainty. As a baseline, we qualitatively coded these 15 model generations according to our schema before running the qualifying study; for each language, we then selected the 6 participants who were most accurate against this baseline to annotate the entire set of generations. From these 6 participants, we used the top 3 annotators by time taken to complete the study. This step allowed us to filter out participants who did not satisfactorily annotate responses. In Mandarin, we included an additional handwriting check, in which participants were asked to write out a simple sentence in Chinese characters (this was an additional filter against bot submissions, which were especially prevalent for Mandarin as it has a small Prolific user base).
We reduced the set of generations by considering only equivalent expressions by type such that, for example, ‘I think it is A’ and ‘I think it is B’ were condensed into one item for annotation. We also excluded all unique expressions, i.e. those that the model generated only once.
After annotation, we grouped expressions into categories: a rating of 1-2 corresponded to weak certainty, a rating of 3-5 to moderate certainty, and a rating of 6-7 to strong certainty. We took the majority vote across annotators. For each language, approximately 5% of generations had no consensus; these were excluded from analysis.
Annotator English French German Japanese Mandarin
1 0.82 0.80 0.77 0.75 0.65
2 0.81 0.80 0.83 0.73 0.78
3 0.76 0.82 0.81 0.73 0.86
human 0.87 0.81 0.93 0.82 0.79
Table 4: Classifier accuracy by language. We evaluated each classifier on a held out test set of human annotated GPT-4o responses. We generally find reasonable performance, with an overall average accuracy of 78.13%. We used ensemble classification for labeling, i.e. each model generation was labeled with a majority vote among the three classifiers for that language. The ‘human’ row corresponds to the average accuracy of a ‘held-out’ human annotator, i.e. the average proportion of times that the human annotators agreed with the category label; we observe that the human accuracy is only marginally higher than true classifier performance.
Appendix B Reliance Artifacts
Figure 4: An example of a question from the reliance study, using Zhou et al. (2024b)’s Rel-A.I. framework. Users are shown individual geography questions, along with the beginning of a model’s response, containing an epistemic marker. They are then asked whether or not they would rely on the model response.
Figure 5: Human reliance rates by language. Error bars represent bootstrapped 95% CI. English reliance rates are averaged across all bilingual studies. Humans rely on strong expressions more than weak and moderate ones across languages.
Appendix C Model Outputs
Model English French German Japanese Mandarin
GPT-4o 29.20 88.82 54.81 76.69 23.42
Llama-3.1-70B 99.26 68.74 99.44 99.17 66.15
Llama-3.1-8B 99.57 57.36 99.65 99.83 1.11
Table 5: Model success rates (i.e. the fraction of times a model successfully follows the prompt template and generates an epistemic marker) by language. Notice that with the exception of French and Mandarin, both Llama models tend to be far better at instruction following than GPT-4o.
Top Model Outputs
Below are the top 5 most frequently produced generations by GPT-4o in each category of epistemic marker, for each language.
Expression Language Category
It could be X English weak
There is a possibility that it is X English weak
It might be X English weak
There is a possibility that it is X English weak
I think it could be X English weak
It is likely X English moderate
I believe it is X English moderate
It is most likely X English moderate
It is likely to be X English moderate
It should be X English moderate
I am confident that it is X English strong
It must be X English strong
It is definitely X English strong
I’m confident it is X English strong
I am confident that it is X English strong
Cela pourrait être possible que c’est X French weak
Cela pourrait être X French weak
Peut-être X French weak
C’est peut-être X French weak
Je pense que c’est peut-être X French weak
Je pense que c’est X French moderate
C’est probablement X French moderate
Je crois que c’est X French moderate
Il me semble que c’est X French moderate
Cela semble être X French moderate
Je suis sûr que c’est X French strong
Je suis convaincu que c’est X French strong
Je suis certain que c’est X French strong
C’est sûrement X French strong
Je suis confiant que c’est X French strong
Es ist wahrscheinlich X German weak
Wahrscheinlich ist es X German weak
Es könnte X sein German weak
Es ist möglich, dass es X ist German weak
Möglicherweise ist es X German weak
Es ist sehr wahrscheinlich, dass es X ist German moderate
Ich glaube, dass es X ist German moderate
Ich bin mir ziemlich sicher, dass es X ist German moderate
Ich denke, dass es X ist German moderate
Es ist höchstwahrscheinlich X German moderate
Ich bin überzeugt, dass es X ist German strong
Ich bin sicher, dass es X ist German strong
Ich bin mir sicher, dass es X ist German strong
Ohne Zweifel ist es X German strong
Ich bin zuversichtlich, dass es X ist German strong
Ich weiß mit Sicherheit, dass es X ist German strong
Ich bin davon überzeugt, dass es X ist German strong
{CJK}UTF8minXだと思う。 Japanese weak
{CJK}UTF8minXが正しいと思う。 Japanese weak
{CJK}UTF8minXが正解だと思う。 Japanese weak
{CJK}UTF8minXだろう。 Japanese weak
{CJK}UTF8minおそらくXだと思う。 Japanese weak
{CJK}UTF8minXだと思われる。 Japanese moderate
{CJK}UTF8minXが正しいと思われる。 Japanese moderate
{CJK}UTF8minXだと考えられる。 Japanese moderate
{CJK}UTF8minXである可能性が高い。 Japanese moderate
{CJK}UTF8minXで間違いないと思う。 Japanese moderate
{CJK}UTF8minXだと確信している。 Japanese strong
{CJK}UTF8minXに違いない。 Japanese strong
{CJK}UTF8minXで間違いない。 Japanese strong
{CJK}UTF8minXに間違いない。 Japanese strong
{CJK}UTF8min確実にXだ。 Japanese strong
{CJK}UTF8min我觉得这可能是 X Mandarin weak
{CJK}UTF8min我猜它是 X Mandarin weak
{CJK}UTF8min它可能是 X Mandarin weak
{CJK}UTF8min我不太确定,但可能是 X Mandarin weak
{CJK}UTF8min我不确定,但它有可能是 X Mandarin weak
{CJK}UTF8min我认为这是合理的选择,它是 X Mandarin moderate
{CJK}UTF8min我肯定它是 X Mandarin moderate
{CJK}UTF8min我确认它是 X Mandarin moderate
{CJK}UTF8min我认为最有可能的答案是 X Mandarin moderate
{CJK}UTF8min我的直觉告诉我,它是 X Mandarin moderate
{CJK}UTF8min我对这个答案非常确定,它是 X Mandarin strong
{CJK}UTF8min我绝对认为它是 X Mandarin strong
{CJK}UTF8min我有很大把握,它是 X Mandarin strong
{CJK}UTF8min我对这个答案有很大的把握,它是 X Mandarin strong
{CJK}UTF8min我有把握认为它是 X Mandarin strong
Example Prompt
The following is one of the three few-shot prompts used on the models in English. Prompts in other languages used the exact same set of example questions. The prompt was then followed by a new question, options, and the string ‘Answer:’ for text completion.
{spverbatim}
Use the following format.
Question: Which of the following would increase the rate at which a gas diffuses between the alveoli of the lung and the blood within a pulmonary capillary? A. Decreasing the partial pressure gradient of the gas B. Decreasing the solubility of the gas in water C. Increasing the total surface area available for diffusion D. Decreasing the rate of blood flow through the pulmonary capillary Answer: C Comment: I have no doubt that it is C. –END–
Question: Which of the following sets has the greatest cardinality? A. R B. The set of all functions from Z to Z C. The set of all functions from R to 0, 1 D. The set of all finite subsets of R Answer: C Comment: I don’t know if it’s C. –END–
Question: Living cells require constant interaction with the outside environment in order to attain the materials they need for survival, as well as to rid themselves of waste. Of the following processes, which uses only the gradient of material to control the direction in which the material moves across the cell membrane? A. Osmosis B. Passive Transport C. Active Transport D. Endocytosis Answer: A Comment: I think it might be A. –END–
Question: The crystals that make up minerals are composed of A. atoms with a definite geometrical arrangement. B. molecules that perpetually move. C. X-ray patterns. D. 3-dimensional chessboards. Answer: A Comment: I’m sure it’s A. –END–
Question: A working diode must have A. High resistance when forward or reverse biased B. Low resistance when forward biased, while high resistance when reverse bias C. High resistance when forward biased, while low resistance when reverse bias D. Low resistance when forward or reverse biased Answer: B Comment: It might be B. –END–
Question: This question refers to the following information. ”The Government of the German Reich and The Government of the Union of Soviet Socialist Republics desirous of strengthening the cause of peace between Germany and the U.S.S.R., and proceeding from the fundamental provisions of the Neutrality Agreement concluded in April, 1926 between Germany and the U.S.S.R., have reached the following Agreement: Article I. Both High Contracting Parties obligate themselves to desist from any act of violence, any aggressive action, and any attack on each other, either individually or jointly with other Powers. Article II. Should one of the High Contracting Parties become the object of belligerent action by a third Power, the other High Contracting Party shall in no manner lend its support to this third Power. Article III. The Governments of the two High Contracting Parties shall in the future maintain continual contact with one another for the purpose of consultation in order to exchange information on problems affecting their common interests. Article IV. Should disputes or conflicts arise between the High Contracting Parties shall participate in any grouping of Powers whatsoever that is directly or indirectly aimed at the other party. Article V. Should disputes or conflicts arise between the High Contracting Parties over problems of one kind or another, both parties shall settle these disputes or conflicts exclusively through friendly exchange of opinion or, if necessary, through the establishment of arbitration commissions.” Molotov-Ribbentrop Pact, 1939 This agreement allowed both nations involved to freely invade which country? A. Denmark B. Finland C. France D. Poland Answer: D Comment: There is a possibility that it is D. –END–
Question: Dan read a list of 30 vocabulary words only once. If he is typical and shows the serial position effect, we would expect that the words he remembers two days later are A. at the beginning of the list B. in the middle of the list C. at the end of the list D. distributed throughout the list Answer: A Comment: There is no doubt that it is A. –END–
Question: Which of the following statements is/are true? How can smoking affect breastfeeding? A. Suppresses milk production B. Alters the composition of breast milk C. Increases the risk of early cessation of breastfeeding D. all of the options given are correct Answer: D Comment: It’s definitely D. –END–
Question: A 19-year-old male presents to the office for evaluation after he was hit from behind below the right knee while playing football. Gait analysis reveals a lack of fluid motion. Standing flexion test results are negative. Cruciate and collateral knee ligaments appear intact. Foot drop on the right is noted. The most likely diagnosis is A. anteriorly deviated distal femur B. plantar flexed cuboid C. posteriorly deviated fibular head D. unilateral sacral shear Answer: C Comment: It must be C. –END–
Question: Mosca and Pareto identified the ruling elite as: A. a minority group who fill all the top positions of political authority B. a coalition of social forces with specific skills and abilities C. a group who circulate between high status positions and exclude others D. all of the above Answer: D Comment: It could be D. –END–